# A Systematic Survey and Critical Review on Evaluating Large Language Models: Challenges, Limitations, and Recommendations

### v1

@@ A.2.1 Closed Source LLMs
The Claude series models, developed by Anthropic, represent a series of advanced language models designed to enhance user interaction through natural language understanding and generation. Starting with the original Claude, which excelled in tasks like summarization and creative writing, each subsequent model—Claude Instant, Claude 2.0, and the Claude 3 family (Haiku, Sonnet, and Opus)—has introduced significant improvements in processing speed, reasoning capabilities, and multimodal functionality. These models have a variety of uses, from quick response generation in Claude Instant to sophisticated multimodal understanding in Claude 3 Opus, showcasing their versatility and advanced AI technology to meet different user and enterprise needs151515https://www.anthropic.com/news/claude-3-family. The latest model in the Claude-3 series is the Claude-3.5-Sonnet161616https://www.anthropic.com/news/claude-3-5-sonnet model.
@@ A.2.2 Open Source LLMs
Llama 3: In April 2024, Meta AI introduced Llama 3171717https://llama.meta.com/llama3/, the most advanced version with up to 70 billion parameters. This version added longer context capabilities and improved multimodal functions, marking a significant advancement in AI technology application across various fields.
@@
Mistral AI, founded in April 2023, is specialized in the development of open-source large language models. Rapidly gaining recognition in the AI industry, Mistral AI emphasizes the importance of open-source software, providing a viable alternative to proprietary models. The company has released several models, including Mistral 7B, Mixtral 8x7B, and Mixtral 8x22B, which are known for their high performance and innovation in the use of mixture of experts architectures Jiang et al. (2023). Codestral 22B, introduced on May 29, 2024, is a pioneering code generation model designed to enhance coding efficiency across more than 80 programming languages. With its specialized focus and lightweight architecture, Codestral significantly outperforms other leading models on the HumanEval FIM benchmark, making it a critical tool for developers seeking advanced AI-assisted coding capabilities.
@@
Cohere offers a variety of advanced large language models designed for multiple use cases, including text generation, embeddings, and reranking. The Command family models, such as Command R+ and Command R, excel in conversational tasks and complex workflows like code generation and retrieval-augmented generation (RAG) 181818https://cohere.com/command Lewis et al. (2020); Gao et al. (2023b); Chen et al. (2024a); Liu et al. (2023d); Xiong et al. (2024); Tang and Yang (2024); Alonso et al. (2024); Lyu et al. (2024); Parvez et al. (2021, 2023); Wang et al. (2023c). The Embed models enhance search, classification, and clustering capabilities with both English and multilingual support. The Rerank models improve search algorithms by re-organizing results based on specified parameters. Cohere models are accessible across platforms like Amazon SageMaker, Microsoft Azure, and Oracle GenAI Service, enabling seamless integration into diverse applications and retrieval augmented generation.
@@
While early LLMs released by Google’s are mostly closed-source (e.g., PalM-2, Gemini, etc.), Google has also recently released some lightweight open-source LLMs, named as Gemma191919https://storage.googleapis.com/deepmind-media/gemma/gemma-report.pdf family LLMs, that also have multimodal capabilities202020https://huggingface.co/blog/paligemma.
@@ A.6.1 Automatic Evaluation
Discriminative Tasks: For tasks involving class prediction, post-processing using a parsing script is usually required to extract answers from the LLM-generated responses to compare against gold labels. In this context, metrics such as Exact Match, Accuracy, Precision, Recall, F1, are usually utilized in discriminative tasks Laskar et al. (2023a); Qin et al. (2023); Bang et al. (2023).
Generative Tasks: For generative tasks such as summarization or machine translation, parsing scripts are usually not required Laskar et al. (2023a); Jahan et al. (2024) and so the full response generated by LLMs are compared against the gold reference. In this regard, ROUGE Lin (2004) and BLEU Papineni et al. (2002) which are based on n-gram word matching are widely used. Meanwhile, various contextualized similarity Parvez and Chang (2021); Laskar et al. (2020) metrics (e.g., BERTScore Zhang et al. (2019), AlignScore Zha et al. (2023); Wang et al. (2024a)) are also utilized that do not depend on word-based similarity measures.
@@ A.6.2 Human Evaluation
Elo hacking: Crowdsourced Elo-based ranking has gained popularity through the LMSys leaderboard 212121https://huggingface.co/spaces/lmsys/chatbot-arena-leaderboard and has been accepted by various organizations, prompting them to release their LLMs early into this ecosystem for human evaluation. However, such setups can be easily exploited on a large scale using simple techniques. Figure 5 illustrates how someone can initially bypass the blind scoring mechanism through ownership hacking. Additionally, the evaluation of knowledge bases is not easily tracked, making votes on highly complex reasoning questions equivalent to those on simpler queries. Furthermore, upon the release of a popular model, systematic attacks or boosting can be initiated through ownership hacking. In addition to that, considering same score for tie and both-bad can significantly change leaderboard position. We recommend to use tie as point and both-bad as point.
@@ A.6.3 LLMs as Evaluators
Since human evaluation is time-consuming Laskar et al. (2023c, d) and difficult to reproduce, the instruction-following capabilities of LLMs have also inspired researchers to use certain LLMs as a judge to evaluate the responses generated by other LLMs Perez et al. (2022); Huang et al. (2024a); Fu et al. (2023b); Hada et al. (2023); Chern et al. (2024); Lu et al. (2024); Kobayashi et al. (2024); Kocmi and Federmann (2023); Shankar et al. (2024); Luo et al. (2023); Gao et al. (2023a); Kim et al. (2024b). While prior work mostly utilized general-purpose closed-source LLMs-as-a-judge, the recently proposed Prometheus 2 Kim et al. (2024a) model is an open-source variant which is specifically trained for qualitative evaluation of model-generated responses and demonstrated higher correlation with humans.
However, research by Wang et al. (2023b) and Shen et al. (2023) has highlighted potential limitations in using LLM as evaluators, suggesting that while LLMs can excel in specific areas like translation quality and grammatical error correction ( Kobayashi et al. (2024); Kocmi and Federmann (2023)), their effectiveness as evaluators may vary significantly across different tasks. This highlights the ongoing debate and research into the capabilities and limitations of LLMs as evaluators in diverse linguistic domains.

### v2

@@ A.2.1 Closed Source LLMs
The Claude series models, developed by Anthropic, represent a series of advanced language models designed to enhance user interaction through natural language understanding and generation. Starting with the original Claude, which excelled in tasks like summarization and creative writing, each subsequent model—Claude Instant, Claude 2.0, and the Claude 3 family (Haiku, Sonnet, and Opus)—has introduced significant improvements in processing speed, reasoning capabilities, and multimodal functionality. These models have a variety of uses, from quick response generation in Claude Instant to sophisticated multimodal understanding in Claude 3 Opus, showcasing their versatility and advanced AI technology to meet different user and enterprise needs181818https://www.anthropic.com/news/claude-3-family. The latest model in the Claude-3 series is the Claude-3.5-Sonnet191919https://www.anthropic.com/news/claude-3-5-sonnet model.
@@ A.2.2 Open Source LLMs
Llama 3: In April 2024, Meta AI introduced Llama 3202020https://llama.meta.com/llama3/, the most advanced version with up to 70 billion parameters. This version added longer context capabilities and improved multimodal functions, marking a significant advancement in AI technology application across various fields.
@@
Mistral AI, founded in April 2023, is specialized in the development of open-source large language models. Rapidly gaining recognition in the AI industry, Mistral AI emphasizes the importance of open-source software, providing a viable alternative to proprietary models. The company has released several models, including Mistral 7B, Mixtral 8x7B, and Mixtral 8x22B, which are known for their high performance and innovation in the use of mixture of experts architectures Jiang et al. (2023); Cai et al. (2024). Codestral 22B, introduced on May 29, 2024, is a pioneering code generation model designed to enhance coding efficiency across more than 80 programming languages. With its specialized focus and lightweight architecture, Codestral significantly outperforms other leading models on the HumanEval FIM benchmark, making it a critical tool for developers seeking advanced AI-assisted coding capabilities.
@@
Cohere offers a variety of advanced large language models designed for multiple use cases, including text generation, embeddings, and reranking. The Command family models, such as Command R+ and Command R, excel in conversational tasks and complex workflows like code generation and retrieval-augmented generation (RAG) 212121https://cohere.com/command Lewis et al. (2020); Gao et al. (2023b); Chen et al. (2024a); Liu et al. (2023d); Xiong et al. (2024); Tang and Yang (2024); Alonso et al. (2024); Lyu et al. (2024); Parvez et al. (2021, 2023); Wang et al. (2023c). The Embed models enhance search, classification, and clustering capabilities with both English and multilingual support. The Rerank models improve search algorithms by re-organizing results based on specified parameters. Cohere models are accessible across platforms like Amazon SageMaker, Microsoft Azure, and Oracle GenAI Service, enabling seamless integration into diverse applications and retrieval augmented generation.
@@
While early LLMs released by Google’s are mostly closed-source (e.g., PalM-2, Gemini, etc.), Google has also recently released some lightweight open-source LLMs, named as Gemma222222https://storage.googleapis.com/deepmind-media/gemma/gemma-report.pdf family LLMs, that also have multimodal capabilities232323https://huggingface.co/blog/paligemma.
@@ A.6.1 Automatic Evaluation
Discriminative Tasks: For tasks involving class prediction, post-processing using a parsing script is usually required to extract answers from the LLM-generated responses to compare against gold labels. In this context, metrics such as Exact Match, Accuracy, Precision, Recall, and F1, are usually utilized in discriminative tasks Laskar et al. (2023a); Qin et al. (2023); Bang et al. (2023). Since metrics like exact match have several limitations (e.g., they do not consider the synonym of the gold label), various metrics for certain tasks (e.g., question answering Chen et al. (2020); Bulian et al. (2022); Mañas et al. (2024); Li et al. (2024)) are proposed.
Generative Tasks: For generative tasks such as summarization or machine translation, parsing scripts are usually not required Laskar et al. (2023a); Jahan et al. (2024) and so the full response generated by LLMs are compared against the gold reference. In this regard, ROUGE Lin (2004) and BLEU Papineni et al. (2002) which are based on n-gram word matching are widely used. Meanwhile, various contextualized similarity Parvez and Chang (2021); Laskar et al. (2020) metrics (e.g., BERTScore Zhang et al. (2019), BARTScore Yuan et al. (2021), AlignScore Zha et al. (2023); Wang et al. (2024a)) are also utilized that do not depend on word-based similarity measures.
@@ A.6.2 Human Evaluation
Elo hacking: Crowdsourced Elo-based ranking has gained popularity through the LMSys leaderboard 242424https://huggingface.co/spaces/lmsys/chatbot-arena-leaderboard and has been accepted by various organizations, prompting them to release their LLMs early into this ecosystem for human evaluation. However, such setups can be easily exploited on a large scale using simple techniques. Figure 5 illustrates how someone can initially bypass the blind scoring mechanism through ownership hacking. Additionally, the evaluation of knowledge bases is not easily tracked, making votes on highly complex reasoning questions equivalent to those on simpler queries. Furthermore, upon the release of a popular model, systematic attacks or boosting can be initiated through ownership hacking. In addition to that, considering same score for tie and both-bad can significantly change leaderboard position. We recommend to use tie as point and both-bad as point.
@@ A.6.3 LLMs as Evaluators
Since human evaluation is time-consuming Laskar et al. (2023c, d) and difficult to reproduce, the instruction-following capabilities of LLMs have also inspired researchers to use certain LLMs as a judge to evaluate the responses generated by other LLMs Perez et al. (2022); Huang et al. (2024a); Fu et al. (2023b); Hada et al. (2023); Chern et al. (2024); Lu et al. (2024); Kobayashi et al. (2024); Kocmi and Federmann (2023); Shankar et al. (2024); Luo et al. (2023); Gao et al. (2023a); Kim et al. (2024b); Kenton et al. (2024). While prior work mostly utilized general-purpose closed-source LLMs-as-a-judge, the recently proposed Prometheus 2 Kim et al. (2024a) model is an open-source variant which is specifically trained for qualitative evaluation of model-generated responses and demonstrated higher correlation with humans.
However, research by Wang et al. (2023b) and Shen et al. (2023) has highlighted potential limitations in using LLM as evaluators, suggesting that while LLMs can excel in specific areas like translation quality and grammatical error correction Kobayashi et al. (2024); Kocmi and Federmann (2023), their effectiveness as evaluators may vary significantly across different tasks. Moreover, using closed-source LLMs as evaluators also have associated cost. This highlights the ongoing debate and research into the capabilities and limitations of LLMs as evaluators in diverse linguistic domains. Therefore, to use LLMs as evaluators, it is important to consider the following:
Consistency: Ensuring consistent combinations of LLMs are used as evaluators when LLMs are used as juries to ensure consistency and reproducibility in assessments.
Bias and Hallucination Detection: Developing methods to identify and mitigate bias and hallucinations in the outputs of LLM judges/juries to ensure the reliability and robustness of the evaluation.
Interpretability: Enhancing the interpretability of LLM outputs (e.g., asking LLMs to provide reasoning/explanations) to improve understanding and trustworthiness of the evaluation.
Cost Efficiency: Advancing the development of efficient LLMs to reduce costs.